0 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

In general, z-test is used when the sample size is large and standard deviation is known.

T-test is more adaptable than Z-test since Z-test will often require certain conditions to be reliable.

If the population standard deviation (σ)\left( \sigma \right) is known, then z test is used.

If the population standard deviation (σ)\left( \sigma \right) is unknown, then t test is used.

Fundamentals

Test statistic for one sample t-test is,

t=xˉμsnt = \frac{{\bar x - \mu }}{{\frac{s}{{\sqrt n }}}}

The formula of degrees of freedom for tt - test is,

df=n1df = n - 1

Here, nn is sample size.

Decision rule:

1.If P-value is less than the Level of significance then reject the null hypothesis, otherwise fail to reject the null hypothesis.

2.If t-test statistic value is greater than the t tabulated value then reject the null hypothesis otherwise fail to reject the null hypothesis.

(a)

From the given information, n=15,t=1.66,α=0.05n = 15,t = 1.66,\alpha = 0.05 .

The degrees of freedom is,

df=n1=151=14\begin{array}{c}\\df = n - 1\\\\ = 15 - 1\\\\ = 14\\\end{array}

The critical value is,

tα/2=TINV(0.05,14)(Usetheexcelfunction)=2.145\begin{array}{c}\\{t_{\alpha /2}} = {\rm{TINV(0}}{\rm{.05,14)}}\left( {{\rm{Use the excel function}}} \right)\\\\ = 2.145\\\end{array}

Since the test statistic value is less than the critical value, do not reject the null hypothesis.

(b)

From the given information, n=15,t=1.66,α=0.05n = 15,t = - 1.66,\alpha = 0.05 .

The degrees of freedom is,

df=n1=151=14\begin{array}{c}\\df = n - 1\\\\ = 15 - 1\\\\ = 14\\\end{array}

The critical value is,

tα/2=T.INV(0.05/2,14)(Usetheexcelfunction)=2.145\begin{array}{c}\\{t_{\alpha /2}} = {\rm{T}}{\rm{.INV(0}}{\rm{.05/2,14)}}\left( {{\rm{Use the excel function}}} \right)\\\\ = - 2.145\\\end{array}

Since the test statistic value does not lie in the rejection region, do not reject the null hypothesis.

(c)

From the given information, n=26,t=2.55,α=0.01n = 26,t = - 2.55,\alpha = 0.01 .

The degrees of freedom is,

df=n1=261=25\begin{array}{c}\\df = n - 1\\\\ = 26 - 1\\\\ = 25\\\end{array}

The critical value is,

tα/2=T.INV(0.01/2,25)(Usetheexcelfunction)=2.787\begin{array}{c}\\{t_{\alpha /2}} = {\rm{T}}{\rm{.INV(0}}{\rm{.01/2,25)}}\left( {{\rm{Use the excel function}}} \right)\\\\ = - 2.787\\\end{array}

Since the test statistic value does not lie in the rejection region, do not reject the null hypothesis.

(d)

From the given information, n=26,t=3.95,α=0.01n = 26,t = - 3.95,\alpha = 0.01 .

The degrees of freedom is,

df=n1=261=25\begin{array}{c}\\df = n - 1\\\\ = 26 - 1\\\\ = 25\\\end{array}

The critical value is,

tα/2=T.INV(0.01/2,25)(Usetheexcelfunction)=2.787\begin{array}{c}\\{t_{\alpha /2}} = {\rm{T}}{\rm{.INV(0}}{\rm{.01/2,25)}}\left( {{\rm{Use the excel function}}} \right)\\\\ = - 2.787\\\end{array}

Since the test statistic value lies in the rejection region, reject the null hypothesis.

Ans: Part a

Do not reject the null hypothesis. There is not sufficient evidence that the true diameter differs from 0.5 in.

Part b

Do not reject the null hypothesis. There is not sufficient evidence that the true diameter differs from 0.5 in.

Part c

Do not reject the null hypothesis. There is not sufficient evidence that the true diameter differs from 0.5 in.

Part d

Reject the null hypothesis. There is sufficient evidence that the true diameter differs from 0.5 in.